OpenStudy (anonymous):

how do you cancel a 3/2 power?

OpenStudy (anonymous):

can u give an example...not quite sure what u meant

OpenStudy (wolf1728):

You don't just cancel a 3/2 power. What exactly are you trying to do?

OpenStudy (anonymous):

like how you would use a square root to cancel something squared

OpenStudy (anonymous):

\[v ^{3/2} for instance\]

OpenStudy (wolf1728):

Well just to make sure you know what a 3/2 power is, 2^(3/2) = sq root (2^3)

OpenStudy (anonymous):

to me its the same as this:\[\sqrt{v^{3}} \]

OpenStudy (wolf1728):

yes that's right

OpenStudy (anonymous):

thats what i was looking for/ i dont think its going to do what i had intentions for though

OpenStudy (wolf1728):

You are probably looking for something like this right? 2^(3/2) = sq root (8) squaring both sides 8 = 8

OpenStudy (wolf1728):

[x^(3/2)]^2 = x^3

OpenStudy (anonymous):

\[v _{y}^{3/2}=mg/b(1-e ^{(-b/m)t})\]

OpenStudy (anonymous):

i was trying to get vy by itself

OpenStudy (wolf1728):

wow - that's a bit difficult. You could square both sides and you'd have vy^3 = [mg/b(1−e(−b/m)t)]^2 which might make an even bigger mess.

OpenStudy (anonymous):

yeah thats what i was thinking. ill probably leave it the way it is. i had to find a equation of the terminal speed under a drag force that was proprtional to v\[_{y ^{3/2}}\]

OpenStudy (agent0smith):

Put both sides to the power of the reciprocal of the power. \[\huge v^{\frac{ 3 }{ 2 }} = x\] put both sides to the power 2/3 \[\huge (v^{\frac{ 3 }{ 2 }})^{\frac{ 2 }{ 3 }} = x ^{\frac{ 2 }{ 3 }}\] which gives \[\huge v = x ^{\frac{ 2 }{ 3 }}\]

OpenStudy (agent0smith):

That's how you get rid of any power - put both sides to the reciprocal of the power.

OpenStudy (anonymous):

thank you sir, thats exactly what i was looking for

OpenStudy (wolf1728):

4(3/2) = 8 So, 4 = 8^(2/3) and it does Good work agent smith

OpenStudy (agent0smith):

:) well, that's the same thing you're doing when you take the square root of both sides of an equation - putting both sides to the power of 1/2, to get rid of a power of 2.

OpenStudy (anonymous):

thats a good explanation